OpenStudy (anonymous):

3x-6y=12 2x+6y=-12 Solve by using elimination

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what are your thoughts?

OpenStudy (anonymous):

i have no idea im so confused

OpenStudy (anonymous):

i forgot how to do it

OpenStudy (wolf1728):

Pandabear78693 - so you were taught how to do these?

OpenStudy (amistre64):

elimination involves adding the equations together so that one of the terms zeros out ...

OpenStudy (anonymous):

yes i was taught to do this but i forgot how to do it @wolf1728

OpenStudy (amistre64):

do the equations have a similar term to them that will cancel when added up?

OpenStudy (anonymous):

3x-6y=12 + 2x+6y=-12 _________ 5x=0 x=0 (Since the y's already canceled each other, I didn't have to multiply anything.) Now plug the x-value in both equation and see if it works: 3(0)-6y=12 -6y=12 y=-2 2(0)+6y=-12 6y=-12 y=-2 So the answer is (0,-2) I think Hope this helps!

OpenStudy (anonymous):

so i will add 3x and 2x

OpenStudy (amistre64):

you will add all like terms, the idea is to get one of the terms to cancel itself out i the process.

OpenStudy (amistre64):

recall that a-a = 0; known as the inverse of addition property or some such

OpenStudy (anonymous):

ooo okk

OpenStudy (anonymous):

thx 4 da medal

OpenStudy (anonymous):

Thanks for helping @amistre64 and @Peaches15 btw ur welcome

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

thank you thanks for explaining it @amistre64